OpenStudy (anonymous):

y varies directly as x. When y = 15, x = 4. Find the constant of variation and use it to determine the value for y when x = 6.

OpenStudy (anonymous):

set it up as a proportion.

OpenStudy (anonymous):

Then, wouldn't you do cross mulitply & divide ?

OpenStudy (anonymous):



OpenStudy (anonymous):

y=k*x

OpenStudy (anonymous):

k is a constant that vary

OpenStudy (anonymous):

15=k4, k=4/15, then use it in the y=4/15*6

OpenStudy (softballgirl372015):

k=15/4

OpenStudy (softballgirl372015):

then you do 15/4 * 6

OpenStudy (anonymous):

yep, sorry wrong calculation